In 15 years, A would be twice as old as he was 10 years ago. How old is he now?
Answer by ikleyn(53751)

.
In 15 years, A would be twice as old as he was 10 years ago. How old is he now?
----------------------------------------------------------

x + 15 = 2(x-10)

is your equation.

Solve it:

x + 15 = 2x - 20,

x = 15 + 20 = 30.

Answer. 35 years.